OpenStudy (anonymous):

5 Graphing Systems of Inequalities questions??? HELP please!!!

OpenStudy (zehanz):

What is your question?

OpenStudy (zehanz):

Name the system you have to graph...

OpenStudy (anonymous):

The graph below represents which system of inequalities?

OpenStudy (anonymous):

y <= -2x + 3 y <= x + 3 y>= -2x + 3 y >= x + 3 y <= -3x + 2 y <= -x + 2 y > -2x + 3 y > x + 3

OpenStudy (anonymous):

Thats question 1

OpenStudy (zehanz):

First, you need to find the equations of the straight lines, in the form y=ax+b. What is the equation of the line through (0,2) and (2, 0)?

OpenStudy (anonymous):

Uhhhhh y=2x+0?

OpenStudy (anonymous):

@ZeHanz

OpenStudy (zehanz):

y=ax+b. b is the y-coordinate of the intersection with the y-axis: b = 2. a is the slope. Slope is -1, because, if you're on the line, and you do one step to the right, you go 1 down. So it is y = -x+2. If you look at the options, there is already only one left...

OpenStudy (anonymous):

The third option?

OpenStudy (anonymous):

i kinda understand how to do it now @ZeHanz